Given the PDF for X, find P( X > 5 ) and find h.

The PDF in uniform with lowest value = 3 and highest value = 10.
Answer #1

Given the PDF for *, find plx>5) f find h The pdf in uniform with lowest value =3& value 10. highest uniform X re b= 10 (a-?wit = (5-01) = 7 I oslz dae Ļ = 91 S=2 dep rajta op fins it S (s<xd mon 10 + 10-3 ㅜ T fx(x) =n= ol 3 fie) (s<x)d